Hi! I would be so glad if anyone can help me on solving this distance-based problem. The "10 hours" thing at the end of the question just left me confused at that point. It is as follows:
An automobile leaves City A for City B 2 hours later than a truck. How many miles per hour must the automobile travel to reach City B at the same time as the truck if the truck travels at an average of 30 miles per hour and makes the trip in 10 hours?

Answer by mananth(16949)

Let us first analyse the truck movement
speed = 30 mph time taken = 10 hours so distance between A & B = 300 miles
The truck has traveled 2 hours when the automobile starts
So it has traveled 60 miles in the 2 hours.
so remaining distance = 240 miles to city B
Let the speed of the automobile = x
time taken to reach the city by the automobile = 300/x
But the truck takes 8 hours to reach the distance at the same time

300/x = 8
8x=300
x=300/8
x= 37.5 mph which is the speed of the automobile